OpenStudy (anonymous):

4. y – 3 = –7 (1 point) 5. x + 4 = 8 (1 point) 6. 7t = 42 (1 point) 7. m ÷ 8 = –4 (1 point) 8. –90 = 10f (1 point)

OpenStudy (anonymous):

Solve each equation or inequality.

OpenStudy (anonymous):

4. y=-4 5. x=4 6. t=42/7=6 8. f=-90/10=-9

OpenStudy (anonymous):

7. m/8=-4 m=-4*8 m= -32